OpenStudy (anonymous):

The vertices of a parallelogram are A(x1, y1), B(x2, y2), C(x3, y3), and D(x4, y4). Which of the following must be true if parallelogram ABCD is proven to be a rectangle? WILL METAL AND FAN!!!!

rebeccaxhawaii (rebeccaxhawaii):



rebeccaxhawaii (rebeccaxhawaii):

those are the points so i dont know what your question is asking

OpenStudy (anonymous):

MY FIRST CHOICE IS A= (Y4-Y-3/X4-X3= Y2-Y3/X3-X2) AND Y4-Y3/X4-X3 TIMES Y3-Y2/X3-X2

OpenStudy (anonymous):

OH AND THAT = -1

OpenStudy (anonymous):

LOOKING FOR AN ANSWER LIKES THAT

OpenStudy (anonymous):